North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Home Locomotives Q6 63395 (NER T2 2238) Q6 Blog Wednesday 17th April 2019

Wednesday 17th April 2019

A bit of a cool overcast start to the day at Grosmont, but warming up with the sun shining by lunchtime. With Steve Hyman, Nigel Bill, Bryan Orange, Nigel Hall, Ian Pearson and myself. Ian lit the stove and made a cuppa (the 'right' colour this time - I have been sacked!) and Nigel brought a lovely fruit cake.

Steve set about working on the right hand rear steam chest cylinder cock, with Ian assisting where necessary. After some discussion about this cylinder cock's carrot valve, and whether it should be given a skim, Ian took it to the machine shop to get some advice from Owain. He thought we should have a new one made, but he couldn’t do one at the moment as he had a long list of priority work because of the necessary repair work following the recent accident involving Eustace Forth and Eric Treacy. Owain suggested that when the T2 is in for its next washout it could be removed and remade. So the cylinder cock and carrot valve were re-fitted. All the cylinder cock linkage bolts were double nutted where no split pin holes were found, and those with holes were split pinned. When Ian was talking to Owain he mentioned the bush for the way shaft pin. Owain said he will come up and measure the hole but it would have to wait until his work load had eased. Nigel H rubbed down the back of the tender which he had painted on Tuesday, then gave it a coat of varnish. I spent the morning washing up, necessitating a visit to the Co-op for a scouring sponge, and topping up the water bottles.

Bryan and Nigel B were busy discussing all the safety aspects of working for NELPG in Deviation Shed and also other locations. One of the outcomes of an ongoing safety review is that there is now a separate signing on point in Deviation Shed itself - it is on the right hand door of the double doors that lead to the stairs in the workshop. Please remember to sign in and out there, but also that you must continue to sign in and out at the MPD signing on point as normal, so the duty fitter knows who is in and out at the end of his day.

Lunch was had around the stove. After lunch, quite a lot of of ash had accumulated in the pit so was removed by Nigel B. Subsequently, I cleaned out the drainage channel and removed the contents. I had been down to the MPD to see Barney and Piglet. Because of the dire shortage of motive power at the moment, they asked if the T2 could be pressed into traffic early if the situation became desparate. After discussion with Nigel and Ian, and today with Bill, we have agreed that the T2 could be made available next week if absolutely necessary, provided the rocking grate and the way shaft pin are sorted, and subject to prior agreement. In any case, a steam test is planned for next week, in preparation for the 1 May centenary event. Black (for the tender sides) and red (for the front buffer beam) plastic will be required if it is to do any running. As a contingency, Ian and Nigel B oiled up the axlebox tops. Last job yesterday, after Nigel H had finished varnishing the back of the tender, was to dismantle the scaffolding and return it all back to the MPD boiler shop. This was done by all our team.

Lots of visitors in Deviation Shed today but Bryan and I did the PR bit and gave them lots of information about NELPG. As a result, we think we might have recruited one working volunteer member and a couple of potential JVs. The new safety valve castings for the T2 and J27 were collected from Lanes in Middlesbrough by Sue Smeaton, and are now at Grosmont ready for machining. On the Railway, trains were being worked with No 926 Repton (after some leaking tubes had been dealt with in the morning), B1 No 1264 (with steam leaking from the left hand cylinder) and J27 No 65894. A certain John Hunt was driving, and he reported the J27 was going well considering it was hauling 6 coaches, and the afternoon FTR exam reported no problems. It has been running since Sunday and should be running all week and over Easter